OpenStudy (anonymous):

Find the slope of a line perpendicular to y=-2/5x+4/5 5/4 5/2 2/5 -5/2

OpenStudy (scarlettred):

is this multiple choice?

OpenStudy (anonymous):

yes maam

OpenStudy (scarlettred):

than -5/2 would be the answer

OpenStudy (anonymous):

mahalo

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

@scarlettred

OpenStudy (anonymous):

actually no. It's supposed to be the opposite

OpenStudy (jdoe0001):

what's the slope of -> y=-2/5x+4/5 ?

OpenStudy (anonymous):

what is it then ? @jdoe0001 @melody16

OpenStudy (jdoe0001):

well.. what do you think is the slope of y=-2/5x+4/5 ?

OpenStudy (anonymous):

If you know what the slope of that equation is then it's the opposite.

OpenStudy (anonymous):

2/5 ? @melody16

OpenStudy (anonymous):

idk, i suck at math badly @jdoe0001

OpenStudy (jdoe0001):

hmm well. notice your equation is in slope-intercept form, that means \(\large \begin{array}{llll} y=&{\color{brown}{ \frac{2}{5}}}x&+\frac{4}{5}\\ &\uparrow &\uparrow \\ &{\color{brown}{ slope}}&y-intercept \end{array}\)

OpenStudy (jdoe0001):

see the slope now?

OpenStudy (scarlettred):

but theres the negative sign in front of the 2/5

OpenStudy (anonymous):

it's a negative slope j

OpenStudy (jdoe0001):

ohh yeah... shoot ok \(\large \begin{array}{llll} y=&{\color{brown}{ -\cfrac{2}{5}}}x&+\cfrac{4}{5}\\ &\uparrow &\uparrow \\ &{\color{brown}{ slope}}&y-intercept \end{array}\)

OpenStudy (jdoe0001):

see the slope of that equation now?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

so now we know the slope is -2/5 the slope of a line that is PERPENDICULAR to that equation line, will be NEGATIVE RECIPROCAL that is \(\bf -\cfrac{2}{{\color{brown}{ 5}}}\qquad negative \to +\cfrac{2}{{\color{brown}{ 5}}}\qquad reciprocal\to +\cfrac{{\color{brown}{ 5}}}{2}\)